1. In a clinical study, 4200 subjects were vaccinated with a flu vaccine manufactured by growing cells in fertilized chicken eggs. Over a period of roughly 28 weeks, 32 of these subjects developed the flu.

You must do all of your calculations on the TI calculator. DO NOT Use the Stats -> Tests for any of the problems.

a. Find the observed statistic. Use the proper notation ( x or p ˆ ). [2pts]

b. Verify that the sample size is large enough to use the normal distribution to construct a confidence interval for the proportion of the population that were vaccinated with the vaccine but still developed the flu. [2 pts]

c. Construct a 95% confidence interval for the proportion of the population that were vaccinated with the vaccine but still developed the flu. [5pts]

d. Provide an interpretation of your interval in the context of the data. [2pts]

e. According to your confidence interval in c), is it reasonable to assume that fewer than 1% of all people vaccinated will develop the flu? Clearly answer “yes” or “no”. _______. And provide an explanation for your answer. [2pts]
Answer #1

Solution :

a) For the given scenario the most appropriate statisic will be sample proportion. The sample proportion of subjects who were vaccinated with the vaccine but develoed the flu is given by,

\large \hat{p} = \frac{32}{4200} = 0.0076

The observed value of the statistic is p̂ = 0.0076.

b) We can use normal distribution to construct a confidence interval if np̂ ≥ 10 and n(1 - p̂) ≥ 10.

We have, n = 4200 and  p̂ = 0.0076

np̂ = 31 which is greater than 10.

n(1 - p̂) = 4168 which is greater than 10.

Hence, sample size is large enough to use the normal distribution.

c) The 95% confidence interval for population proportion is given as follows :

\large \left ( \hat{p} \pm Z_{\frac{0.05}{2}}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}} \right )

Where, p̂ is sample proportion, n is sample size and Z(0.05/2) is critical z-value to construct 95% confidence interval.

We have,  p̂ = 0.0076, n = 4200

Using the calculator we get, Z(0.05/2) = 1.96

Hence, the 95% confidence interval for the proportion of the population that were vaccinated with the vaccine but still developed the flu is,

\large \left (0.0076 \pm 1.96\times\sqrt{\frac{0.0076\times(1-0.0076)}{4200}} \right )

\large =\left (0.0076 \pm 0.0026 \right )

\large =\left (0.0076 - 0.0026, 0.0076 + 0.0026 \right )

\large =\left (0.0050, 0.0102 \right )

The 95% confidence interval for the proportion of the population that were vaccinated with the vaccine but still developed the flu is (0.0050, 0.0102).

d) Interpretation : We are 95% confident that the true proportion of the population that were vaccinated with the vaccine but still developed the flu lies between 0.0050 and 0.0102.

e) Since, the 95% confidence interval also contain the value greater than 1% = 0.01, there it is not reasonable to assume that fewer than 1% of all people vaccinated will develop the flu.